Fill in the P(X=×) value in the table below to give a legitimate probability distribution for the discrete random variable X, whose possibles values are.
-1, 0, 1, 4
Value x of X.                    P(X=×)
-1.                                     0.16
0.                                     0.15
1                                      0.28
4.                                      _____
5.                                      _____
Answer #1

Ans:

For a valid probability distribution,sum of probabilities should be equal to 1.

P(x=-1)+P(x=0)+P(x=1)+P(x=4)=1

0.16+0.15+0.28+P(x=4)=1

P(x=4)=1-0.16-0.15-0.28

P(x=4)=1-0.59

P(x=4)=0.41